# A framework for the use of generative modelling in non-equilibrium statistical mechanics

### v3

@@ 3 Some preliminary philosophical considerations
How might one apply the FEP to a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
@@
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. Even if one could, the internal states would be forever sequestered behind blanket states. Applications of the FEP place the generative model centre-stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
@@ 4 A map of that part of the territory that behaves as if it were a map
5 Some remarks on terminological idiosyncrasies of the FEP literature
In summary, we have argued that the map-territory fallacy, as it has been leveraged against the FEP (e.g., by [41, 42]) does not apply to FEP-theoretic modelling—that is, not in the sense that it conflates the model and the thing modelled. We claim this allegation constitutes a fallacy (one might call this a map-territory fallacy fallacy).
Mathematically, there is no ambiguity or conflation of map and territory in FEP-theoretic modelling. Although the language used to narrate the mathematics of the FEP are at times ambiguous (which may have led to the confusion in philosophical commentary on the FEP), the mathematical structure itself is not. In the use of the FEP, an object is modelled as deploying a variational density, which its internal states parameterise; its dynamics are given in our generative model, which is meant to capture or represent the real (or most likely) dependence relations of the real-world system [8]. It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
Charitably, one might argue that the allegations that FEP-theoretic modelling commits the map-territory fallacy may be due to several terminological ambiguities and idiosyncrasies that have mired the FEP literature. In fairness, these ambiguities and idiosyncrasies lend themselves to misinterpretation. We now discuss two such cases.
5.1 Generative models and variational densities versus recognition densities
The first concerns the ambiguous discussion or description of two core constructs in the FEP literature—namely, the generative model and the variational (or recognition) density. It is noteworthy that the manner in which these constructs are deployed in the FEP-theoretic formulation is markedly different from previous Bayesian approaches. This has seldom been noted in the literature (with a few exceptions, e.g., [8]).
In more traditional machine learning approaches, e.g., in the literature on the Helmholtz machine [52], the generative model and the recognition model correspond to the feedforward and the feedback passes that together constitute the process of data-driven Bayesian inference. In these settings, both the generative model and the recognition model are straightforwardly encoded in the connection weights of the Helmholtz machine: the generative model is encoded in the weighting of the backwards connections, which establish an information channel from high-level latent spaces of the neural network to its sensory layers; whereas the recognition model is encoded in the feedforward connections. In this setting, one assumes that any system that implements this pair of functionally differentiated weightings (e.g., a brain) implements a generative and recognition model pair.
Much of the literature on the FEP is based on this reading of the generative and recognition models: it is often said within this literature that FEP-theoretic generative models are encoded by the brain, much like these would be in a Helmholtz machine. However, whilst generative and variational models are also used in the FEP-theoretic formulation, they are deployed in a markedly different way, which we discussed extensively above.
Overlooking these differences may have contributed to the impression that theorists of the FEP conflate the metaphorical map and the territory. Recall that a generative model, generically, is just a joint probability density defined over the states (or paths) of a system. In other words, it is our map or representation of the system being modelled and its mechanics; and more particularly, it is our model of the dependencies between the variables of the system, providing us with equations of motion for the whole system (i.e., it is defined over internal, blanket, and external states/paths).
Technically, we say that the system that we are modelling entails a generative model; whereas the internal states of the particle (look, to the modeller, as if they) estimate the parameters of a conditional density over external states. This difference is crucial to understanding what is at stake in FEP-theoretic modelling: we say that one subset of variables over which the generative model is defined (i.e., the internal states) encodes a variational or recognition model over another subset (the external states). Entailment is the property that, metaphorically, “the part of the territory that behaves like a map” (the internal states that encode the parameters of the variational density) is itself part of the territory described by our map (the generative model) and so is subject to the mechanics of that territory, as described by our map. Entailment means that the relationships that hold between variables, as we represent in the generative model, are obtained as a description of the dynamics of the real-world system.
The subtle but key point is that the generative model is not encoded by the internal states of the particle being considered: it is not some kind of internal model. Rather, the generative model is a description of the dependence relations that obtain in the dynamics of a system: i.e., such-and-such dynamics entail such-and-such equations of motion and dependencies. FEP-theoretic modelling does not assume that the generative model is encoded in the brain. Arguably, to claim this would indeed constitute a map-territory fallacy. Under the FEP, parts of the system itself (i.e., by internal states) look to an external observer as if, or can be modelled as if, they encode the parameters of a variational model.
However, this raises a question: if it is our scientific model, how can the generative model play a role in inference if only a conditional posterior model is encoded by the system? To claim that a system uses or leverages its generative model is, in some sense, a shorthand expressing the fact that the system leverages its own causal structure, and the data patterns that it elicits through action, to generate patterns of adaptive behaviour—all of which are subject to the regularities that are captured in the generative model. This is a statement of embodied and enactive cognition [53, 8]. The system, in having sensory states, also has access to the causal structure in the generative model.
So, the generative model is our model or map, but it denotes the causal structure of the system being modelled (i.e., the territory). This causal structure—and in particular, the presence of a Markov blanket or particular partition—means that some parts of the system (the internal states) encode a variational density over some other parts of the system (the external states); that is, they “act like a map” or model the statistics of external states.
5.2 Generative models and generative processes
The second ambiguity concerns the generative model and the so-called “generative process”. Some presentations of the FEP sometimes define the generative model as a scientific model of some data-generating generative process; see, for instance, [54, 8]). On this reading, the generative model is our “map” of some real-world process, namely, the generative process—which would then correspond to the metaphorical territory.
In practice, however, in FEP-theoretic modelling, this distinction does not map onto that between a scientific model and real-world system, i.e., onto the distinction between the map and territory. Perhaps confusingly, the generative process does not denote (the structure of) the real data-generating process that we are trying to model. In FEP-theoretic formulations, the generative process is also part of our scientific model.
It should be noted that, practically, the distinction between generative model and generative process only comes into play when we are implementing the dynamics of a specific system in simulations. When we write code to simulate the system of interest, typically, we need to specify (i) the dependencies that hold between the variables of a system, which we call the generative model; and (ii) the equations of motion that govern the time evolution of the external states per se in our simulation. The latter is labelled the ‘generative process’, since it generates the dynamics of external states, which are tracked by internal states. The key thing to note, however, is that generative process is also part of our scientific model—it harnesses what we know about the physics of the situation described by the generative model, in which the thing is embedded and ‘lives’.

### v4

@@ 4 Some preliminary philosophical considerations
Throughout, ‘as if’ marks an interpretive equivalence: internal flows that minimise variational free energy can be read as variational inference without presupposing explicit, propositional models inside the system. Where specific mechanisms are identified (e.g., precision-weighted message passing), this reading coincides with literal implementation; otherwise, it remains an explanatory stance. How might one apply the FEP to make a model of a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
@@
In the context of this meta-question, one may ask how an observer could come to consciously think an observed system is manipulating a model if the observer’s own brain is simply flowing down a free-energy gradient. Here, an observer’s conscious report that ‘X is using an internal model’ may arise from some higher-order generative model over another agent’s (or one’s own) internal states, i.e., beliefs about the other system’s blanket-to-internal mappings (what we have called ‘observing the observer’). Conscious thought in this context is a higher-level subset of those internal states which summarises beliefs about others (and the self) within a hierarchical model. Mechanistically, this report is implemented by the observer’s own free-energy-minimising dynamics.
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. Even if one could, the internal states would be forever sequestered behind blanket states. Applications of the FEP place the generative model centre-stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket. Here both observed and observer are described at the mechanistic level by flows on free energy flows; at the informational level, the observer’s higher-order model licenses attributions like ‘the observed system is manipulating an internal model’.